# Relation-Aware Diffusion Model for Controllable Poster Layout Generation

### v1

@@ Abstract.
$\ast$ Both authors contributed equally to this research
† Work done during an internship at JD.com
@@ 1. Introduction
The generation of poster layouts is geared towards forecasting the location and category of visual elements on an image. This task is crucial for both the aesthetic appeal and information dissemination of posters. Owing to the requirement of simultaneously accounting for graphic interrelationships and image compositions to create top-notch poster layouts, this demanding task is usually completed by professional designers. However, manual design is time-consuming and financially burdensome.
@@ 2.1. Layout Generation
According to whether the visual content is considered, we divide the deep generative models into two categories: content-agnostic and content-methods. Content-agnostic methods usually yield layouts with visual balance and symmetry as there are fewer constraints, making them suitable for documents, user interfaces, and publication generation. LayoutVAE (Jyothi et al., 2019), which utilizes Variational Autoencoders, is a method that learns to produce layouts based on the categories of elements. To further improve the quality of the generated layouts, transformers (Kong et al., 2022; Yang et al., 2021) are used in the generation task. Due to the attention mechanism, transformer-based methods are capable of implicitly learning the relationships between elements.
@@ 2.2. Diffusion Models
In recent years, diffusion models (Song et al., 2020; Ho et al., 2020) have gradually become the focus of generative tasks because of their impressive high-quality generative capabilities. The diffusion and denoising processes are key components of this approach. Diffusion refers to the gradual transformation of an initial image into a final noisy image through a series of small, random perturbations. Denoising, on the other hand, is the process of learning to remove noise from the image to actual distribution. Besides image generation, Diffusion models are gaining momentum in various fields and show promising performance. InST (Zhang et al., 2023) implemented Inversion-Based Style Transfer with Diffusion Models. Video LDM (Blattmann et al., 2023) achieved high-resolution video generation by training a diffusion model in a compressed low-dimensional latent space. Naturally, the diffusion model is also introduced into the field of layout generation. LayoutDM (Inoue et al., 2023) used a discrete diffusion model to predict the attributes of elements like category and position. LDGM (Hui et al., 2023) unifies unconditional and conditional generation in a single diffusion model. But these methods are oblivious to input contents and perform poorly in poster layout generation. By introducing a multimodal diffusion model, our method can align the image and texts and produce more visually convincing posters.
@@ 4.2.2. Text Encoder
Given all the promotional slogans of the product on a poster, we extract textual features through a pre-trained language model RoBERTa (Liu et al., 2019). We note that the product description is not simply repeating the product name, but highlighting the selling points of the product. For instance, if you want to promote a computer, you describe it as ”high CPU performance” without mentioning ”computer”. Therefore, it is important to narrow the gap between the product description and the product itself. To address the problem, we gathered a vast product corpus of 200 million items from a Chinese e-commerce platform and adapt the same pretraining strategy which comprises Masked Language Model (MLM), Attribute-Value Prediction (AVP), and Tertiary Category Prediction (TCP) to finetune RoBERTa. For MLM, we randomly mask certain words from the input product title and feed it into the language model. This allows the model to predict the original sentence accurately. AVP and TCP are used to predict the value of a product based on its attribute and tertiary category. AVP is utilized to extract product values from the product description by utilizing product attribute queries. TCP involves the analysis and assessment of product information to determine the appropriate category. In order to let the model perceive the relationship between text length and layout, we supplement textual length embedding as a part of text features. Finally, we fuse the content features and length features of the text by concat operation, as the output of the text encoder, denoted as $L\in\mathbb{R}^{D_{n}\times d}$. It is worth noting that our method is not limited to Chinese. Migrating to another language only requires replacing the text encoder here.
@@ 5.3. Comparison with Content-Aware Methods
As mentioned in the previous chapters, ContentGAN and CGL-GAN are two generators considering the influence of image content on layout, so here is our main comparison model. We re-implement ContenGAN based on the released codes [^1] https://xtqiao.com/projects/content aware layout, and specifically add content feature extraction and text feature extraction modules consistent with our method. Meanwhile, we tried our best to re-implement the CGL-GAN method based on the details in the paper. The quantitative comparison results of the three methods are shown in Tab.  1. No matter whether in user study or composition-relevant metric, our method is obviously winning, which shows that the proposed method has a better ability to represent the relationship between image content and layout.
@@ 5.5. Controllable Layout Generation
In order to verify whether the output results of the model are acceptable given the part layout, we conduct different experiments and the results are shown in Fig.  10. Our model can give qualified results, especially in the results of the third column, our model will not generate additional layouts without enough layout space, which shows that the model has strong constraints and generalization ability.
@@ 5.6. Ablation Studies
Table 3. Ablation studies of VTRAM. $Ours^{*}$ means our model without VTRAM.
@@
Table 4. Ablation studies of GRAM. $Ours^{*}$ means our model without GRAM.

### v2

@@ Abstract.
copyright: acmcopyrightjournalyear: 2023copyright: acmlicensedconference: Proceedings of the 32nd ACM International Conference on Information and Knowledge Management; October 21–25, 2023; Birmingham, United Kingdombooktitle: Proceedings of the 32nd ACM International Conference on Information and Knowledge Management (CIKM ’23), October 21–25, 2023, Birmingham, United Kingdomprice: 15.00doi: 10.1145/3583780.3615028isbn: 979-8-4007-0124-5/23/10ccs: Computing methodologies Neural networks
@@ 1. Introduction
Poster layout generation aims to predict the position and category of graphic elements on the image, which is important for visual aesthetics and information transmission of posters. Due to the need to consider both graphic relationships and image compositions when creating high-quality poster layouts, this challenging task is usually completed by professional designers. However, manual design is often time-consuming and financially burdensome.
@@ 2.1. Layout Generation
According to whether the visual content is considered, we divide the deep generative models into two categories: content-agnostic and content-aware methods. Content-agnostic methods usually yield layouts with visual balance and symmetry as there are fewer constraints, making them suitable for documents, user interfaces, and publication generation. LayoutVAE (Jyothi et al., 2019), which utilizes Variational Autoencoders, is a method that learns to produce layouts based on the categories of elements. To further improve the quality of the generated layouts, transformers (Kong et al., 2022; Yang et al., 2021) are used in the generation task. Due to the attention mechanism, transformer-based methods are capable of implicitly learning the relationships between elements.
@@ 2.2. Diffusion Models
In recent years, diffusion models (Song et al., 2020; Ho et al., 2020) have gradually become the focus of generative tasks because of their impressive high-quality generative capabilities. The diffusion and denoising processes are key components of this approach. Diffusion refers to the gradual transformation of an initial image into a final noisy image through a series of small, random perturbations. Denoising, on the other hand, is the process of learning to remove noise from the image to actual distribution. Besides image generation, Diffusion models are gaining momentum in various fields and showing promising performance. DiffusionDet (Chen et al., 2022) is the first to apply diffusion model for the task of object detection. InST (Zhang et al., 2023) implemented Inversion-Based Style Transfer with Diffusion Models. Video LDM (Blattmann et al., 2023) achieved high-resolution video generation by training a diffusion model in a compressed low-dimensional latent space. Naturally, the diffusion model is also introduced into the field of layout generation. LayoutDM (Inoue et al., 2023) uses a discrete diffusion model to predict the attributes of elements like category and position. LDGM (Hui et al., 2023) unifies unconditional and conditional generation in a single diffusion model. But these methods are oblivious to input contents and perform poorly in poster layout generation. By introducing a multimodal diffusion model, our method can align the image and texts and produce more visually convincing posters.
@@ 4.2.2. Text Encoder
Given all the promotional slogans of the product on a poster, we extract textual features through a pre-trained language model RoBERTa (Liu et al., 2019). We note that the product description is not simply repeating the product name, but highlighting the selling points of the product. For instance, if you want to promote a computer, you describe it as ”high CPU performance” without mentioning ”computer”. Therefore, it is important to narrow the gap between the product description and the product itself. To address the problem, we gathered a vast product corpus of 200 million items from JD.com and adapt the same pretraining strategy which comprises Masked Language Model (MLM), Attribute-Value Prediction (AVP), and Tertiary Category Prediction (TCP) to finetune RoBERTa. For MLM, we randomly mask certain words from the input product title and feed it into the language model. This allows the model to predict the original sentence accurately. AVP and TCP are used to predict the value of a product based on its attribute and tertiary category. AVP is utilized to extract product values from the product description by utilizing product attribute queries. TCP involves the analysis and assessment of product information to determine the appropriate category. In order to let the model perceive the relationship between text length and layout, we supplement textual length embedding as a part of text features. Finally, we fuse the content features and length features of the text by concat operation, as the output of the text encoder, denoted as $L\in\mathbb{R}^{D_{n}\times d}$. It is worth noting that our method is not limited to Chinese. Migrating to another language only requires replacing the text encoder here.
@@ 5.3. Comparison with Content-Aware Methods
As mentioned in the previous chapters, ContentGAN and CGL-GAN are two generators considering the influence of image content on layout, so here is our main comparison model. We re-implement ContentGAN based on the released codes [^1] https://xtqiao.com/projects/content aware layout, and specifically add content feature extraction and text feature extraction modules consistent with our method. Meanwhile, we tried our best to re-implement the CGL-GAN method based on the details in the paper. The quantitative comparison results of the three methods are shown in Tab.  1. No matter whether in user study or composition-relevant metric, our method is obviously winning, which shows that the proposed method has a better ability to represent the relationship between image content and layout.
@@ 5.5. Controllable Layout Generation
Given partial layout. In order to verify whether the output results of the model are acceptable given the part layout, we conduct different experiments and the results are shown in Fig.  10. Our model can give qualified results, especially in the results of the third column, our model will not generate additional layouts without enough layout space, which shows that the model has strong constraints and generalization ability.
@@ 5.6. Ablation Studies
Table 3. Ablation studies of VTRAM. $*{}^{*}$ means our model without VTRAM.
@@
Table 4. Ablation studies of GRAM. $*{}^{*}$ means our model without GRAM.